OpenStudy (anonymous):

Some determine for me the roots x^4 +6x^2 -27=0

OpenStudy (anonymous):

make substitution x^2=y then the eqn. is (y+9)(y-3)=0 which gives x=+rt(3),-rt(3),3i,-3i

OpenStudy (anonymous):

\[x^4+6x^2-27=(x^2+9)(x^2-3)=0\]

OpenStudy (anonymous):

are you working with complex numbers or just real? if real the answers are \[\pm \sqrt{3}\]

OpenStudy (anonymous):

real no.s

OpenStudy (anonymous):

oops sorry saubhik got it already

OpenStudy (anonymous):

oh well in real numbers \[x^2+9=0\] has no solution because if you write \[x^2=-9\] you are looking for a real number whose square is negative and there isn't one.